• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name lkaufman

    Log entry time 20:14:09 on June12,2004

    Entry number 125471

    This entry is a followup to: 125467

    keyword=Wrong left arm optics during owl.

    Because the left dipole was out of tolerance for the momentum setting we were running with, (2.87 rather than 3.009), I spent some time getting the magnet to the correct setting by putting in the current manually (although it turned out the quadrupoles were set incorrectly).  I now understand after talking to Javier, the reason that the dipole would not come up to the correct setpoint.
    Please follow instructions in halog #125467 if the dipole is ever reset!
    Note: The left arm optics were incorrect for all runs during the owl shift and the beginning of day shift - at least HAPPEX DAQ Runs 2043-2049 are affected.  The RHWP scans, are fine, runs 2046-2048, because they don't use data from detectors downstream of the target.